The distance between the cathode (filament) and the target in an X-ray tube is 1.5 m. If the cutoff wavelength is 30 pm, find the electric field between the cathode wavelength decrease?

Grade:11

1 Answers

Aditi Chauhan
9 years ago
d = 1.5 m, λ = 30 pm = 30 * 10^-3 nm E = hc/λ = 1242/30* 10^3 = 41.4 * 10^eV Electric field = V/d = 41.4* 10^3/1.5 = 27.6 * 10^3 V/m = 27.6 KV/m.